OpenStudy (anonymous):

Find all the real square roots of -1/-121 a) -1/11 b) 1/11, -1/11 c) 11, -11 d) none

OpenStudy (anonymous):

square roots produce positive answers, by definition. So a, b, and c are wrong, but d is also wrong, because (-1)/(-121) has a square root of 1/11. A different question of solve x^2 = (-1)/(-121) has the answer of b.

Parth (parthkohli):

Why not write (-1)/(-121) as 1/121?

OpenStudy (anonymous):

@Parth - I only did it that way to preserve her original question

Parth (parthkohli):

\[\large \rm \sqrt{1 \over 121} = {\sqrt{1} \over \sqrt{121}}\]

OpenStudy (anonymous):

Thanks everyone i think explaining how to get the answer helps me to better understand :) Thanks!